A uniform rod of length 2l is placed with one end in contact with the horizontal table and is then inclined at an angle α to horizontal and allowed to fall.When it becomes horizontal ,it's angular velocity will be 

Grade:11

1 Answers

Vikas TU
5 years ago
From energy conservation, when it is inclined,
U.E = K.E
mg((lsinα)/2) = 0.5*I*w^2
mglsinα = Iw^2
where I is the moment of inertia about the centre of the rod.
I = m(2l)^2/12 => ml^2/3
mglsinα = ((ml^2)*w)/3
gsinα = lw/3
w = (3gsinα)/l rad/s